# Towards Understanding the Mechanisms of Classifier-Free Guidance

### v1

@@ 4 Investigating CFG in Nonlinear Models
We now explore how the findings from the linear analysis extend to real-world diffusion models. Recent studies [31, 21, 22] show that diffusion models transition from a linear regime to a nonlinear regime as the noise level decreases. In the linear regime, where $\sigma(t)$ is large, the learned diffusion denoisers $\mathcal{D}_{\bm{\theta}}$ can be well approximated by the optimal linear denoiser $\mathcal{D}_{\mathrm{L}}$ (7) (see both qualitative and quantitative verification in section F.1). As $\sigma(t)$ decreases, the diffusion model enters the nonlinear regime where $\mathcal{D}_{\bm{\theta}}$ diverges from $\mathcal{D}_{\mathrm{L}}$. Interestingly, this linear-to-nonlinear transition correlates with the coarse-to-fine effects of CFG. As shown in Figures 5 and 26, in the linear regime, linear and nonlinear CFG produce similar effects, substantially reshaping the global structure of the samples. In contrast, in the nonlinear regime, nonlinear CFG primarily refines local details while preserving the overall structure, leading to different effects as those of linear CFG. This linear-to-nonlinear, coarse-to-fine transition motivates our separate analyses of CFG behavior in each regime.
Figure 5: Linear-to-nonlinear transition in diffusion models. (a) and (b) compare nonlinear CFG and linear CFG applied to a deep diffusion model (EDM). The leftmost column shows unguided samples; subsequent columns show final samples when guidance is applied only at a specific noise level, with $\gamma=15$ (See Figure 26 for more examples).
@@ 4.1 CFG in the Linear Regime
Figures 7 and 6 illustrates the effects of separately applying (i) nonlinear CFG, (ii) linear CFG, (iii) mean-shift guidance, (iv) positive CPC guidance, and (v) negative CPC guidance within the linear regime of EDM over a broad range of $\gamma$. As expected, linear CFG [^2] Note that the “linear CFG” here differs from the “linear CFG” in section 3, where both the naive conditional score and the cfg guidance are linear. In contrast, the linear guidance in this section, along with its components, is applied to a real-world deep diffusion model. produces results that closely match those of nonlinear CFG, both significantly altering the overall structures of unguided samples. Notably, decomposing linear CFG provides further insights: Mean-shift guidance dominates CFG in the linear regime. As shown in Figure 7, qualitatively, mean-shift guidance alone replicates the effects of both linear and nonlinear CFG. Consistent with this observation, FDDINOv2 scores confirm that the mean-shift term is the main contributor to CFG’s overall behavior. Because mean-shift term is independent of the sampling trajectory, it reduces sample diversity. As shown in Figure 6, mean-shift guidance improves generation quality only within a limited range of $\gamma$, after which further increases in $\gamma$ degrade FDDINOv2 scores, reflecting a loss of diversity.
@@
Moreover, the observation that the sample-independent mean-shift guidance alone leads to improved FDDINOv2 score implies that simply initializing the sampling process from a mean-shifted Gaussian, $\bm{x}_{T}\sim\mathcal{N}(\gamma(\bm{\mu}_{c}-\bm{\mu}_{uc}),\sigma^{2}(T)\bm{I})$, with no additional guidance applied, can improve the generation quality, which is verified in section F.3. Practically, this initialization trick avoids the per-sample, per-timestep network inference required by CFG, hence could be beneficial in applications where inference speed is important. Theoretically, the observation that the mean-shifted initialization yields better sample quality compared to the commonly used zero-mean initialization suggests the existence of class-specific clusters in the intermediate noisy distribution $p(\bm{x};\sigma(t))$. Understanding how these clusters are formed and what additional structures they possess is an interesting future direction.
CPC guidance also improves generation quality. Although overshadowed by the mean-shift term, applying CPC guidance independently offers notable benefits as well. Qualitatively, positive and negative CPC terms preserve the global structure of unguided samples while refining existing features, remaining effective over a broader range of $\gamma$. Moreover, CPC guidance sometimes mitigate the artifacts introduced by the mean-shift term, such as color oversaturation in the golden retriever example at $\gamma=20$. Lastly, we note that the effects of CPC guidance can vary by class. As shown in Figure 6, negative CPC term improves FDDINOv2 scores for "golden retriever" but has minimal effect on "sports car". These findings are verified on 10 classes, with additional results presented in section F.2, Figures 28, 29, 30 and 31.
Figure 7: Effects of CFG in Linear Regime. Each row demonstrates the impact of different guidance types applied to EDM within the linear regime, with varying guidance strength $\gamma$. The guidance is applied only within intervals specified in the subtitles, where the model exhibits linear behavior. For additional experimental results, please refer to appendix F.
@@ 4.2 CFG in the Nonlinear Regime
In the nonlinear regime where $\sigma(t)$ is small, as shown in Figure 5(b), the effects of linear CFG diverge from those of the actual nonlinear CFG. By Tweedie’s formula, the CFG guidance (5) can be expressed as $g(\bm{x},t)=\frac{\mathcal{D}(\bm{x};\sigma(t),\bm{c})-\mathcal{D}(\bm{x};\sigma(t)}{\sigma^{2}(t)}$, where $\mathcal{D}(\bm{x};\sigma(t),\bm{c})$ and $\mathcal{D}(\bm{x};\sigma(t))$ denote the optimal conditional and unconditional denoisers minimizing  (3). Unlike the linear setting, these denoisers do not admit closed-form expressions in the nonlinear regime, making analytical study challenging. Nevertheless, when denoisers are parameterized with deep networks with no additive ’bias’ parameters, their input-output mappings are locally piecewise linear [32, 33], satisfying:
@@
where $\nabla_{\bm{x}}\mathcal{D}(\bm{x};\sigma(t),\bm{c})$ and $\nabla_{\bm{x}}\mathcal{D}(\bm{x};\sigma(t))$ are the local Jacobians of the denoisers. In this case, CFG guidance becomes $\frac{(\nabla_{\bm{x}}\mathcal{D}(\bm{x};\sigma(t),\bm{c})-\nabla_{\bm{x}}\mathcal{D}(\bm{x};\sigma(t)))\bm{x}}{\sigma^{2}(t)}$ , which shares a similar form as $g_{cpc}(\bm{x},t)$ defined under the linear setting in (12) since $\tilde{\bm{\Sigma}}_{c,t}-\tilde{\bm{\Sigma}}_{uc,t}=\nabla_{\bm{x}}\mathcal{D}_{\mathrm{L}}(\bm{x};\sigma(t),\bm{c})-\nabla_{\bm{x}}\mathcal{D}_{\mathrm{L}}(\bm{x};\sigma(t))$, where $\mathcal{D}_{\mathrm{L}}$ is the optimal linear denoiser. Thus, the guidance can be decomposed into positive and negative CPC components, enhancing the former and suppressing the latter. The key difference from the linear setting is that the CPCs here are adaptive to $\bm{x}$.
The bias-free denoisers fall under the class of pseudo-linear denoisers [34, 35], which admit the form $\mathcal{D}(\bm{x};\sigma(t))=\bm{W}(\bm{x};\sigma(t))\bm{x}$ with $\bm{W}(\bm{x};\sigma(t))$ beging symmetric and input-dependent. Importantly, it is shown in [34] that if the origin is a stationary point of the log-density, i.e., $\nabla_{\bm{x}}\log p(\bm{x};\sigma(t))|_{\bm{x}=\bm{0}}=\bm{0}$, then the optimal denoiser must admit this piecewise linear structure. Even if the diffusion models are not bias-free and the locally linear property does not hold, (16) still serves as a reasonable proxy. As discussed in section 2.3, the Jacobian $\nabla_{\bm{x}}\mathcal{D}(\bm{x};\sigma(t))$ is proportional to the posterior covariance. Its leading singular vectors capture the dominant structures shared by all plausible clean images corresponding to the noisy input $\bm{x}$, while those associated with near-zero singular values span a null space that is irrelevant to the image content. Thus, (16) acts as a weighted projection onto the subspace that captures the most informative image structures.
@@
As shown in Figure 8(b)–(c), both conditional and unconditional Jacobians effectively denoise the input, though their outputs are brighter and sharper than those of the actual denoisers (d)–(e). Comparing  Figure 8(b) and (c), we observe that the conditional and unconditional Jacobians yield denoised outputs with highly similar global structure. However, the conditional Jacobian preserves finer local details. A similar pattern holds for the actual denoisers shown in Figure 8(d) and (e). This suggests that constructing effective guidance requires distinguishing directions that encode class-specific details from those capturing generic structures. Empirically, in Figure 9, the following guidance, inspired by the positive CPC guidance (13), leads to similar effects as the actual CFG:
@@
where $\hat{\lambda}_{+,i}$ and $\bm{v}_{+,i}$ denote the positive eigenvalues and eigenvectors of $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})-\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t))$. Unlike (13), this guidance applies projection to the denoiser’s output rather than the noisy input $\bm{x}_{t}$, which we find leads to better qualitative results. For comparison, we also test the following guidance which enhances all the conditional posterior PCs:
@@
where $\lambda_{c,i}$ and $\bm{u}_{c,i}$ are the singular values and vectors of $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})$. As shown in Figure 9, this approach fails to improve image quality and frequently produces images with oversaturated colors, indicating that not all of these PCs encode the class-specific features.
Note that we do not claim the nonlinear CFG is equivalent to (17), but the qualitative similarity suggests a shared underlying mechanism: identifying and amplifying class-specific features while suppressing generic ones. Due to the black-box nature of deep networks, analyzing the exact mechanism remains difficult, but it is an important direction for future research.
Figure 9: Effects of CFG in the Nonlinear Regime. Different guidance methods, each with a fixed strength of $\gamma=15$, are applied at individual timesteps in the nonlinear regime. Each image shows the final output when guidance is applied solely at the timestep indicated at the top. Note that (17) matches the effects of CFG by enhancing finer image details, whereas (18) does not improve generation quality. For additional experimental results, please refer to Figure 37.
@@ 5 Discussion and Conclusion
Although our analysis is based on linear diffusion model (Gaussian data) assumption, the results remain noteworthy since: (i) CFG significantly enhances the generation quality of linear diffusion models, making the linear setting a meaningful stand-alone testbed for studying CFG and (ii) real-world diffusion models can be well-approximated by their linear counterparts for a wide range of noise levels. We note that the dynamics of linear setting is by itself complex: an interpretable solution to the linear reverse ODE is unattainable unless additional assumptions are imposed on the covariance structures (e.g., the common principal components assumption). A natural next step is to extend the analysis to Gaussian mixtures. We have made some initial attempts in appendix G, showing that CFG guidance in the Gaussian mixture setting can be decomposed in a similar manner as the single Gaussian case.
More broadly, beyond CFG, PCA has been widely utilized for extracting visual features or semantic concepts from diffusion models [36, 37, 38]. Our results indicate that extending these approaches with Contrastive PCA can be a promising next step for more controllable and interpretable generation.
@@ Acknowledgment
We acknowledge funding support from NSF CCF-2212066, NSF CCF- 2212326, NSF IIS 2402950, and ONR N000142512339. We also acknowledge the computing support from NCSA Delta GPU [39]. We thank Mr. Yixiang Dai for fruitful discussions and valuable feedbacks.
@@ B.1 Naive Diffusion Reverse ODE
Let $\bm{\mu}$ and $\bm{\Sigma}$ be the mean and covariance of $p_{\text{data}}(\bm{x})$ respectively. Suppose $\bm{U}\bm{\Lambda}\bm{U}^{T}$ is the full SVD of $\bm{\Sigma}$ with $\bm{U}\in\mathbb{R}^{d\times d}$ being orthonormal and $\bm{\Lambda}=\text{diag}(\lambda_{1},\cdots,\lambda_{d})$ contains the singular values. For image datasets, the covariance is often low-rank implying some singular values are 00. Under the constraint that $\mathcal{D}(\bm{x};\sigma(t))$ is linear (with a bias term), the optimal solution to (3) has the closed-form [21]:

### v3

@@ 4 Investigating CFG in Nonlinear Models
We now explore how the findings from the linear analysis extend to real-world diffusion models. Recent studies [32, 21, 22] show that diffusion models transition from a linear regime to a nonlinear regime as the noise level decreases. In the linear regime, where $\sigma(t)$ is large, the learned diffusion denoisers $\mathcal{D}_{\bm{\theta}}$ can be well approximated by the optimal linear denoiser $\mathcal{D}_{\mathrm{L}}$ (7) (see both qualitative and quantitative verification in section˜F.1). As $\sigma(t)$ decreases, the diffusion model enters the nonlinear regime where $\mathcal{D}_{\bm{\theta}}$ diverges from $\mathcal{D}_{\mathrm{L}}$. Interestingly, this linear-to-nonlinear transition correlates with the coarse-to-fine effects of CFG. As shown in Figures˜5 and 26, in the linear regime, linear and nonlinear CFG produce similar effects, substantially reshaping the global structure of the samples. In contrast, in the nonlinear regime, nonlinear CFG primarily refines local details while preserving the overall structure, leading to different effects as those of linear CFG. This linear-to-nonlinear, coarse-to-fine transition motivates our separate analyses of CFG behavior in each regime.
Figure 5: Linear-to-nonlinear transition in diffusion models. (a) and (b) compare nonlinear CFG and linear CFG applied to a deep diffusion model (EDM). The leftmost column shows unguided samples; subsequent columns show final samples when guidance is applied only at a specific noise level, with $\gamma=15$ (See Figure˜26 for more examples).
@@ 4.1 CFG in the Linear Regime
Figures˜7 and 6 illustrates the effects of separately applying (i) nonlinear CFG, (ii) linear CFG, (iii) mean-shift guidance, (iv) positive CPC guidance, and (v) negative CPC guidance within the linear regime of EDM over a broad range of $\gamma$. As expected, linear CFG [^2] Note that the “linear CFG” here differs from the “linear CFG” in section 3, where both the naive conditional score and the cfg guidance are linear. In contrast, the linear guidance in this section, along with its components, is applied to a real-world deep diffusion model. produces results that closely match those of nonlinear CFG, both significantly altering the overall structures of unguided samples. Notably, decomposing linear CFG provides further insights:
Mean-shift guidance dominates CFG in the linear regime. As shown in Figure˜7, qualitatively, mean-shift guidance alone replicates the effects of both linear and nonlinear CFG. Consistent with this observation, FDDINOv2 scores confirm that the mean-shift term is the main contributor to CFG’s overall behavior. Because mean-shift term is independent of the sampling trajectory, it reduces sample diversity. As shown in Figure˜6, mean-shift guidance improves generation quality only within a limited range of $\gamma$, after which further increases in $\gamma$ degrade FDDINOv2 scores, reflecting a loss of diversity.
@@
Moreover, the observation that the sample-independent mean-shift guidance alone leads to improved FDDINOv2 score implies that simply initializing the sampling process from a mean-shifted Gaussian, $\bm{x}_{T}\sim\mathcal{N}(\gamma(\bm{\mu}_{c}-\bm{\mu}_{uc}),\sigma^{2}(T)\bm{I})$, with no additional guidance applied, can improve the generation quality, which is verified in section˜F.3. Practically, this initialization trick avoids the per-sample, per-timestep network inference required by CFG, hence could be beneficial in applications where inference speed is important. Theoretically, the observation that the mean-shifted initialization yields better sample quality compared to the commonly used zero-mean initialization suggests the existence of class-specific clusters in the intermediate noisy distribution $p(\bm{x};\sigma(t))$. Understanding how these clusters are formed and what additional structures they possess is an interesting future direction.
CPC guidance also improves generation quality. Although overshadowed by the mean-shift term, applying CPC guidance independently offers notable benefits as well. Qualitatively, positive and negative CPC terms preserve the global structure of unguided samples while refining existing features, remaining effective over a broader range of $\gamma$. Moreover, CPC guidance sometimes mitigate the artifacts introduced by the mean-shift term, such as color oversaturation in the golden retriever example at $\gamma=20$. Lastly, we note that the effects of CPC guidance can vary by class. As shown in Figure˜6, negative CPC term improves FDDINOv2 scores for "golden retriever" but has minimal effect on "sports car". These findings are verified on 10 classes, with additional results presented in section˜F.2, Figures˜28, 29, 30 and 31.
Figure 7: Effects of CFG in Linear Regime. Each row demonstrates the impact of different guidance types applied to EDM within the linear regime, with varying guidance strength $\gamma$. The guidance is applied only within intervals specified in the subtitles, where the model exhibits linear behavior. For additional experimental results, please refer to appendix˜F.
@@ 4.2 CFG in the Nonlinear Regime
In the nonlinear regime where $\sigma(t)$ is small, as shown in Figure˜5(b), the effects of linear CFG diverge from those of the actual nonlinear CFG. By Tweedie’s formula, the CFG guidance (5) can be expressed as $g(\bm{x},t)=\frac{\mathcal{D}(\bm{x};\sigma(t),\bm{c})-\mathcal{D}(\bm{x};\sigma(t)}{\sigma^{2}(t)}$, where $\mathcal{D}(\bm{x};\sigma(t),\bm{c})$ and $\mathcal{D}(\bm{x};\sigma(t))$ denote the optimal conditional and unconditional denoisers minimizing  (3). Unlike in the linear setting, these denoisers do not admit closed-form expressions in the nonlinear regime, making analytical study difficult. Nevertheless, when denoisers are parameterized by deep networks with no additive ’bias’ parameters, their input-output mappings are locally piecewise linear [33, 34], satisfying:
@@
where $\nabla_{\bm{x}}\mathcal{D}(\bm{x};\sigma(t),\bm{c})$ and $\nabla_{\bm{x}}\mathcal{D}(\bm{x};\sigma(t))$ are the local Jacobians of the denoisers. In this case, CFG guidance becomes $\frac{(\nabla_{\bm{x}}\mathcal{D}(\bm{x};\sigma(t),\bm{c})-\nabla_{\bm{x}}\mathcal{D}(\bm{x};\sigma(t)))\bm{x}}{\sigma^{2}(t)}$ , which shares a similar form as $g_{cpc}(\bm{x},t)$ defined under the linear setting in (12) since $\tilde{\bm{\Sigma}}_{c,t}-\tilde{\bm{\Sigma}}_{uc,t}=\nabla_{\bm{x}}\mathcal{D}_{\mathrm{L}}(\bm{x};\sigma(t),\bm{c})-\nabla_{\bm{x}}\mathcal{D}_{\mathrm{L}}(\bm{x};\sigma(t))$, where $\mathcal{D}_{\mathrm{L}}$ is the optimal linear denoiser. Thus, the guidance can again be decomposed into positive and negative CPC components, enhancing the former and suppressing the latter. The key distinction from the linear setting is that here, the CPCs are adaptive to $\bm{x}$.
The bias-free denoisers belong to the broader class of pseudo-linear denoisers [35, 36], which admit the form $\mathcal{D}(\bm{x};\sigma(t))=\bm{W}(\bm{x};\sigma(t))\bm{x}$, where $\bm{W}(\bm{x};\sigma(t))$ is symmetric and input-dependent. Importantly, it is shown in [35] that if the origin is a stationary point of the log-density, i.e., $\nabla_{\bm{x}}\log p(\bm{x};\sigma(t))|_{\bm{x}=\bm{0}}=\bm{0}$, then the optimal denoiser must possess such a piecewise linear structure. Even if the diffusion models are not bias-free and the locally linear property does not hold exactly, (16) still serves as a reasonable proxy. As discussed in section˜2.3, the Jacobian $\nabla_{\bm{x}}\mathcal{D}(\bm{x};\sigma(t))$ is proportional to the posterior covariance. Its leading singular vectors capture the dominant structures shared by all plausible clean images corresponding to the noisy input $\bm{x}$, while directions associated with near-zero singular values span a null space irrelevant to the image content. Hence, (16) performs a weighted projection onto the subspace encoding the most informative image structures—effectively functioning as a valid denoising operator.
@@
Indeed, as shown in Figure˜8(b)–(c), both conditional and unconditional Jacobians effectively denoise the input, although their outputs appear brighter and sharper than those from the actual denoisers in (d)–(e). Comparing  Figure˜8(b) and (c), we find that the conditional and unconditional Jacobians yield denoised outputs with similar global structure, which implies both capture the generic structure of the current sample. However, the conditional Jacobian additionally preserves finer, class-specific details. A similar pattern holds for the actual denoisers shown in Figure˜8(d) and (e).
For guidance purposes, our goal is to selectively enhance these fine, class-specific details that the conditional denoiser captures but the unconditional one does not. Achieving this requires identifying directions that encode class-dependent information from those represent generic structures. Empirically, as shown in Figure˜9, the following guidance, inspired by the positive CPC guidance (13), can lead to similar effects as the actual nonlinear CFG, sharpening image details:
@@
where $\hat{\lambda}_{+,i}$ and $\bm{v}_{+,i}$ denote the positive eigenvalues and eigenvectors of $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})-\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t))$. Unlike (13), this guidance applies projection to the denoiser’s output rather than the noisy input $\bm{x}_{t}$, which we find leads to better qualitative results. For comparison, we also test the following non-selective guidance which enhances all the conditional posterior PCs:
@@
where $\lambda_{c,i}$ and $\bm{u}_{c,i}$ are the singular values and vectors of $\nabla\mathcal{D}_{\bm{\theta}}(\bm{x}_{t};\sigma(t),\bm{c})$. As shown in Figure˜9, this approach fails to improve image quality and frequently produces images with oversaturated colors, indicating that not all of these PCs encode the class-specific features—effective guidance must selectively amplify only those that do.
We note that our heuristic guidance serves as a conceptual approximation and may not always perfectly align with actual CFG behavior; in practice, the actual nonlinear CFG yields more stable and consistent results. Due to the black-box nature of deep networks, fully characterizing this mechanism remains challenging, and we regard this as an important direction for future research.
Figure 9: Effects of CFG in the Nonlinear Regime. Different guidance methods, each with a fixed strength of $\gamma=15$, are applied at individual timesteps in the nonlinear regime. Each image shows the final output when guidance is applied solely at the timestep indicated at the top. Note that (17) matches the effects of CFG by enhancing finer image details, whereas (18) does not improve generation quality. For additional experimental results, please refer to Figure˜37.
@@ 5 Discussion and Conclusion
The experiments in the main-text are conducted extensively using the EDM-1 model [4], which operates directly in pixel space with $64\times 64$ resolution. In appendix˜G, we present complementary results on the EDM-2 [5] latent diffusion model, which generates images at $512\times 512$ resolution.
@@
Although our analysis is based on linear diffusion model (Gaussian data) assumption, the results remain noteworthy since: (i) CFG significantly enhances the generation quality of linear diffusion models, making the linear setting a meaningful stand-alone testbed for studying CFG and (ii) real-world diffusion models can be well-approximated by their linear counterparts for a wide range of noise levels. We note that the dynamics of linear setting is by itself complex: an interpretable solution to the linear reverse ODE is unattainable unless additional assumptions are imposed on the covariance structures (e.g., the common principal components assumption). A natural next step is to extend the analysis to Gaussian mixtures. We have made some initial attempts in appendix˜H, showing that CFG guidance in the Gaussian mixture setting can be decomposed in a similar manner as the single Gaussian case.
We believe our findings open several promising directions for future research. First, the observed lack of class-specificity issue implies the current training procedures for diffusion models remain suboptimal. This highlights the need for developing principled training objectives that explicitly encourage the model to learn class-specific patterns. Second, beyond the context of CFG, PCA has been widely utilized for extracting visual features or semantic concepts from diffusion models [37, 38, 39]. Our results suggest that extending these approaches with Contrastive PCA can be a promising next step for more controllable and interpretable generation.
@@ Acknowledgment
We acknowledge funding support from NSF CCF-2212066, NSF CCF- 2212326, NSF IIS 2402950, and ONR N000142512339. This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications[40]. We thank Mr. Yixiang Dai for fruitful discussions and valuable feedbacks.
@@ B.1 Naive Diffusion Reverse ODE
Let $\bm{\mu}$ and $\bm{\Sigma}$ be the mean and covariance of $p_{\text{data}}(\bm{x})$ respectively. Suppose $\bm{U}\bm{\Lambda}\bm{U}^{T}$ is the full SVD of $\bm{\Sigma}$ with $\bm{U}\in\mathbb{R}^{d\times d}$ being orthonormal and $\bm{\Lambda}=\text{diag}(\lambda_{1},\cdots,\lambda_{d})$ contains the singular values. For image datasets, the covariance is often low-rank implying some singular values are 0. Under the constraint that $\mathcal{D}(\bm{x};\sigma(t))$ is linear (with a bias term), the optimal solution to (3) has the closed-form [21]: